OpenStudy (anonymous):

For y = f(x) = x^12 - 12x, and 0≤x≤2, find the value(s) of x for which.... 1. f(x) has a local maximum... (write in ascending order!) x=_______, x=_______ 2. f(x) has a local minimum. If there are none, write NA. x=_______ 3. f(x) has a global maximum. If there are none, write NA. x=_______ 4. f(x) has a global minimum. If there are none, write NA. x=________ Thank you!! so would the first step be to get the derivative and set equal to zero? :/

OpenStudy (unklerhaukus):

take the derivative

OpenStudy (anonymous):

okay:) so \[f'(x)=12x ^{11} - 12\] ?

OpenStudy (unklerhaukus):

a local max or min occurs when f'(x) = 0

OpenStudy (anonymous):

okay so 12(x^11 - 1) = 0 ? is that how i would set it up?

OpenStudy (unklerhaukus):

yeahhhh

OpenStudy (anonymous):

okay so using x^11-1=0 you get x^11= 1 so \[x=\sqrt[11]{1}\] ? is that right? :/

OpenStudy (unklerhaukus):

haha yes

OpenStudy (anonymous):

and also x=0 ? so x= 0 and x= 1 ? :/

OpenStudy (unklerhaukus):

how do you get x=0?

OpenStudy (anonymous):

oh so is it just x=1?

OpenStudy (anonymous):

i wasn't sure, because my problem asks for two local maxes, but i only found x=1 :/ haha

OpenStudy (unklerhaukus):

x=1, is the point , but is is a max or min? is the second derivative (at x=1) positive or negative?

OpenStudy (anonymous):

oh so second derivative is \[f''(x)=132x ^{10}\] ?

OpenStudy (unklerhaukus):

yeah, now is f''(1) positive or negative?

OpenStudy (unklerhaukus):

if it's positive , it mean the graph is curving upwards, ie its a min if negative , curving downwards, ie a max

OpenStudy (anonymous):

positive right? so it's a max?

OpenStudy (unklerhaukus):

um,

OpenStudy (unklerhaukus):

if it's positive , it mean the graph is curving upwards, ie its a min

OpenStudy (anonymous):

ohh so it's a local min? so our local min here is x=1 ?

OpenStudy (unklerhaukus):

yes

OpenStudy (anonymous):

yay! :) so what about the local maxes?

OpenStudy (anonymous):

This graph can give you some ideas plot x^12 - 12 x from x= -1.2 to x=1.4

OpenStudy (anonymous):

See what happens near x=1 http://www.wolframalpha.com/input/?i=plot+x%5E12+-+12+x+from+x%3D+-1.2+to+x%3D-.8

OpenStudy (anonymous):

Sorry near x = -1

OpenStudy (anonymous):

I agree there is only one local minimum at x=1

OpenStudy (anonymous):

okay:) so what about the local maxes? :/ how do we find those?

OpenStudy (anonymous):

This function doe not have a local maximum.

OpenStudy (anonymous):

You need f'(x)=0 and f''(x) < 0 The second derivative of your function is \(132 x^{10} \ge 0 \). So it cannot be negative

OpenStudy (anonymous):

ohh so there are no local maximums? my problem says that f(x) has a local maximum and asks to order them in ascending order... so there are two? confused haha :P

OpenStudy (anonymous):

There is no local maximum everywhere. I did not see that the interval is restricted to \( 0\le x \le 2\). Hold on, we have to consider the end points 0 and 2

OpenStudy (anonymous):

okay:)

OpenStudy (anonymous):

f(0)=0 is a local maximum and f(2)=4072 is also a local maximum

OpenStudy (anonymous):

ohh okay so i would write x=0, x=4072 ?

OpenStudy (anonymous):

or is it x=0 and x=2 ?

OpenStudy (anonymous):

No x=0 and x=2

OpenStudy (anonymous):

ahh okay:) so 1 and 2 are done! what about global max and min?

OpenStudy (anonymous):

At x =0 f has a global minimum at x=2 f has a global maximum

OpenStudy (anonymous):

ohhh so global is still based off what we found from local information?

OpenStudy (anonymous):

x=x0 is a global min means that for every x in [0,2] \[ f(x) \ge f(x0) \]

OpenStudy (anonymous):

x=x0 is a global max means that for every x in [0,2]\[ f(x)\le f(x0)\]

OpenStudy (anonymous):

ahh okay:) awesome!! so this problem is now complete? :O

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

awesome! thank you! :)

OpenStudy (anonymous):

YW